I need to solvethe equation 6/5 +3/2x = 7/15
I have tried multiplying both sides by 30x but end up with x=-45/22 or-2 1/22

Answer by checkley77(12844)

6/5+3/2x=7/15
3x/2=7/15-6/5
3x/2=(7-3*6)/15
3x/2=(7-18)/15
3x/2=-11/15
3x/2=-11/15
15*3x=2*-11
45x=-22
x=-22/45 answer.